# Invisible Walls in Cities: Leveraging Large Language Models to Predict Urban Segregation Experience with Social Media Content

### v3

@@ Abstract
Understanding experienced segregation in urban daily life is crucial for addressing societal inequalities and fostering inclusivity. The abundance of user-generated reviews on social media encapsulates nuanced perceptions and feelings associated with different places, offering rich insights into segregation. However, leveraging this data poses significant challenges due to its vast volume, ambiguity, and confluence of diverse perspectives. To tackle these challenges, we propose using Large Language Models (LLMs) to automate online review mining for segregation prediction. We design a reflective LLM coder to digest social media content into insights consistent with real-world feedback, and eventually produce a codebook capturing key dimensions that signal segregation experience, such as cultural resonance and appeal, accessibility and convenience, and community engagement and local involvement. Guided by the codebook, LLMs can generate both informative review summaries and ratings for segregation prediction. Moreover, we design a REasoning-and-EMbedding (RE’EM) framework, which combines the reasoning and embedding capabilities of language models to integrate multi-channel features for segregation prediction. Experiments on real-world data demonstrate that our framework greatly improves prediction accuracy, with a 22.79% elevation in R2 and a 9.33% reduction in MSE. The derived codebook is generalizable across three different cities, consistently improving prediction accuracy. Moreover, our user study confirms that the codebook-guided summaries provide cognitive gains for human participants in perceiving POIs’ social inclusiveness. Our study marks an important step toward understanding implicit social barriers and inequalities, demonstrating the great potential of promoting social inclusiveness with AI.
1 Introduction
Segregation refers to the systematic separation of individuals or groups based on certain characteristics. Historically, research on segregation has focused primarily on residential patterns Park and Burgess (1925), where individuals of the same racial or ethnic group are more likely to reside in the same neighborhoods Schelling (2006). This type of spatial segregation has been linked to negative social outcomes, such as limited upward social mobility and increased crime rates Gordon and Monastiriotis (2006); Kramer and Hogue (2009). Recent studies have shifted focus to experienced segregation—the dynamic segregation individuals experience in daily movements Moro et al. (2021). Despite the seemingly free human movements in most modern urban spaces, certain demographic groups continue to face barriers to equal access and social interactions. In other words, there seem to be “invisible walls” in cities that prevent sufficient social interactions between groups. These walls are not merely the result of physical proximity but are influenced by cultural, social, and economic factors that drive segregation on a more nuanced level. Accurately predicting segregation experiences can inform individuals to avoid potentially uncomfortable situations and support policymakers in fostering social inclusiveness. Moreover, research suggests that segregation experiences can intensify in larger urban areas Nilforoshan et al. (2023), underscoring the need to explore the underlying mechanisms for growing more equitable cities admist ever-increasing urbanization. However, existing studies only present retrospective and holistic measurements, providing limited insights and prediction capabilities.
@@ 1 Introduction
The rapid advancement of web technologies has enabled fine-grained records of social interactions through diverse media, such as textual posts Weller et al. (2014), photos Gilbert et al. (2013), and short videos Chen et al. (2024). As a result, social media platforms have become extensive social sensing systems, generating rich digital traces that reflect social dynamics in the physical world Stier et al. (2021); Iqbal et al. (2023). Such data offers a unique opportunity to uncover subtle patterns or preferences in individuals’ everyday life, enabling more accurate predictions and insights into how segregation manifests beyond spatial confines. However, identifying cues of segregation experience from massive social media content poses significant challenges, requiring advanced methods to capture nuanced feelings, and there is no established procedure to guide this process.
@@
To address these challenges, we propose to leverage the reasoning power of LLMs for automated social media content analysis. We propose a reflective LLM coder, featuring a strategic agentic workflow that integrates two key components: a reflective attributor and a code summarizer. The reflective attributor employs an abductive reasoning approach: it prompts the LLM to estimate a location’s appeal to different demographic groups from its social media content, and reconcile discrepancies through iterative reflection on observed segregation patterns. Subsequently, the code summarizer applies chain-of-thought reasoning Wei et al. (2022) to iteratively merge insights into a structured codebook. This codebook guides LLMs to transform free-text reviews into structured summaries, highlighting segregation-related factors like cultural resonance and community engagement. Complementing this qualitative approach, we propose a REasoning and EMbedding (RE’EM) framework for quantitative segregation prediction, which combines LLMs’ reasoning capabilities with the representational power of pre-trained embedding models. First, we prompt the LLM to provide structured ratings of a place’s appeal to different groups based on the codebook, allowing the reasoning outcomes of LLMs to be vectorized and easily integrated with other channels of features. Concurrently, we finetune an embedding model to learn global representations optimized for segregation prediction. Finally, we fuse structured ratings, global embeddings, and population information using a neighbor-aware multi-view predictor.
@@
2 Related Work
Experienced Segregation. The study of segregation traces its roots back to the early 20th century, when sociologists examined the division of urban spaces into “ecological niches”, each inhabited by distinct social groups Park and Burgess (1925). Researchers revealed the detrimental effects of enduring segregation even in the absence of “legalized racial segregation”, on education, income, housing, and crime King and Mieszkowski (1973); Massey et al. (1987); Charles (2003). Albeit offering valuable insights, these studies are highly constrained by data availability, overlooking the reality that individuals spend significant time and engage in numerous interactions beyond the confines of their residential neighborhoods Wang et al. (2018). With the rapid proliferation of smart mobile devices, the ability to track people’s intricate movements in urban spaces has emerged Xu et al. (2016). This has paved the way for a novel research avenue aimed at quantifying and understanding the type of segregation individuals actually experience in their daily movements Moro et al. (2021); Athey et al. (2021); de la Prada and Small (2024), including how it may change during disasters Yabe et al. (2023); Chen et al. (2023) and scale with city sizes Nilforoshan et al. (2023). Nevertheless, existing studies primarily focus on static spatial distribution of demographic groups and physical movement, overlooking the complex socioeconomic factors behind the segregation phenomenon, such as cultural resonance and local involvement. Thus, these works can only provide retrospective measurement but have limited predictive power. In contrast, our work establishes an LLM-based method to automatically extract nuanced features from online reviews for experienced segregation prediction.
@@ 2 Related Work
Mining Web Data with LLMs. The digital footprints and rich textual contents people generate on Web platforms have proven to be informative for a wide array of social phenomena, including health outcome Nguyen et al. (2016), mental well-being Mitchell et al. (2013), crime rates Fatehkia et al. (2019), and neighborhood disparities Rama et al. (2020); Iqbal et al. (2023). However, these studies often rely on pre-calculated indices or pre-defined word lists to engineer and extract features. This is not only labor-intensive but also lacks adaptability to evolving research questions and contexts. Given the remarkable language understanding and reasoning capabilities of LLMs, recent research explored the potential of leveraging LLMs to mine Web data for social good, such as revealing food-related social prejudice Luo et al. (2024), evaluating public accessibility Li et al. (2024), and capturing urban perception Santos et al. (2024). Nevertheless, these studies often lack an examined framework for insight extraction from massive social media content, which limits their effectiveness in capturing the complexity of Web data. In contrast, our work presents the first step toward unlocking the reasoning power of LLMs to code free-form online reviews, extracting human-comprehensible, informative insights for segregation prediction.
Figure 1: Demonstration of Reflective Coding.
3 Preliminary
3.1 Problem Formulation
@@
Inspired by Moro et al. (2021), we compute the proportion of visitors at POI $i$ by each racial group $q$, denoted as $\tau_{qi}$. We then quantify the segregation at POI $i$ as the deviation of the visitor proportion $\tau_{qi}$ from the city’s residential proportion $T_{q}$ (k is a constant that normalizes $S_{i}$ to range between 0 and 1):
$S_{i}=k\sum_{q}(|\tau_{iq}-T_{q}|).$ (1)
@@
$min~\text{diff}(f(i,c,p),S_{i}).$ (2)
3.2 Data
@@
Mobility data. To derive the ground truth of experienced segregation, we utilize the Safegraph Patterns Dataset (accessed through Advan) [^3] https://www.deweydata.io/data-partners/advan. This dataset records monthly visits to each POI originating from various CBGs. We aggregate one year’s visitation records to construct a robust dataset with extensive observations. To address the distinct POI indexing systems between Yelp and SafeGraph, we match POIs by ensuring identical names and a location deviation within 200 meters. In Appendix A, we analyze the variability of POI features, highlighting the importance of our study.
4 Reflective LLM Coder
The vast volume of raw social media content creates a significant cognitive barrier to revealing the encoded segregation experience. Therefore, an effective approach is needed to process and extract insights. Building on research demonstrating the consistency and stability of LLM-based coding Dai et al. (2023); Tai et al. (2024), we design a reflective LLM coder to generate an insightful codebook from a small set of POI data, capturing the multi-faceted factors influencing segregation and offering a structured framework for deeper analysis. As shown in Figure 1, it contains two modules, reflective attributor and code summarizer.
4.1 Reflective Attributor
We design a reflective attributor to identify factors influencing a POI’s appeal to different groups. Specifically, we design a two-step Chain-of-Thought (CoT) scheme Wei et al. (2022) that guides an LLM to integrate multiple perspectives from texts and images, forming a reflection-driven process that refines insights based on real-world feedback.
@@ 4.1 Reflective Attributor
Table 1: Automatically constructed codebook.
Reflection-guided insight refinement. Importantly, we introduce real-world data to guide LLM’s reflections upon these insights. By comparing actual visitor demographics with expected residential compositions, the LLM refines the established aspects in an abductive manner, discarding inconsistencies and maintaining alignment with observed behaviors. This step not only enhances the relevance of the insights, but also provides a grounded explanation for observed visitation preferences.
Figure 2: Reasoning-and-Embedding (RE’EM) framework.
@@ 4.2 Code Summarizer
5 REasoning-and-EMbedding (RE’EM) Framework
@@ 5.1 Reasoning Channel
$r_{i},e_{i}=\text{LLM}(u_{i},\text{codebook}).$ (3)
@@
$v_{i}^{r}=\text{ReasonAdapter}(r_{i}).$ (4)
5.2 Embedding Channel
The embedding channel extracts deep semantic representations $v_{i}^{e}$ from the review corpus associated with each POI $i$. Due to the token length constraints, we apply data augmentation, i.e., sample different subsets of $c_{i}$ for input to the embedding model. We fine-tune an open-source text embedding model, GTE-base Li et al. (2023), to enhance its capability of extracting higher-level representations while preserving the pre-trained semantic knowledge. Although both the reasoning and embedding channels take $c_{i}$ as input, they extract complementary aspects: structured attribute reasoning in the former, and holistic semantic embedding in the latter. The resulting embedding is then passed through an embedding adapter (MLP) to produce $v_{i}^{e}$.
$v_{i}^{e}=\text{EmbeddingAdapter}(\text{GTE}(c_{i})).$ (5)
5.3 Population Channel
@@
$v_{i}^{p}=\text{PopulationAdapter}(p_{i}).$ (6)
5.4 Neighbor-Aware Multi-View Predictor
@@
$\hat{S_{i}}=\text{Fuse}(v_{i}^{r},v_{n_{i}}^{r},v_{i}^{e},v_{n_{i}}^{e},v_{i}^{p},v_{n_{i}}^{p}).$ (8)
6 User Study
We conduct a user study to evaluate the cognitive gain to humans provided by our generated codebook. The study involves 75 researchers with hands-on experience or sufficient knowledge in social media studies. To ensure that participants fully understood the survey tasks, we provided clear explanations of the geographic and racial concepts required for the study. Our questionnaire is publicly accessible  [^4] https://drive.google.com/file/d/1-NIjIPiVi_m_ONBFruYYuh4nvyhe4PF-/view.
@@ 6 User Study
Our Model 0.0068±0.0003 9.33%(p=0.0080) 0.0823±0.0021 4.75%(p=0.0071) 0.0662±0.0021 3.07%(p=0.0597) 0.3885±0.0278 22.79%(p=0.0022)
Table 2: Prediction performances.

### v4

# Invisible Walls in Cities: Designing LLM Agent to Predict Urban Segregation Experience with Social Media Content

@@ Abstract.
Understanding experienced segregation in urban daily life is crucial for addressing societal inequalities and fostering inclusivity. The abundance of user-generated reviews on social media encapsulates nuanced perceptions and feelings associated with different places, offering rich insights into segregation. However, leveraging this data poses significant challenges due to its vast volume, ambiguity, and confluence of diverse perspectives. To tackle these challenges, we propose a novel Large Language Model (LLM) agent to automate online review mining for segregation prediction. Specifically, we propose a reflective LLM coder to digest social media content into insights consistent with real-world feedback, and eventually produce a codebook capturing key dimensions that signal segregation experience, such as cultural resonance and appeal, accessibility and convenience, and community engagement and local involvement. Guided by the codebook, LLMs can generate both informative review summaries and ratings for segregation prediction. Moreover, we design a REasoning-and-EMbedding (RE’EM) framework, which combines the reasoning and embedding capabilities of language models to integrate multi-channel features for segregation prediction. Experiments on real-world data demonstrate that our agent substantially improves prediction accuracy, with a 22.79% elevation in R2 and a 9.33% reduction in MSE. The derived codebook is generalizable across three different cities, consistently improving prediction accuracy. Moreover, our user study confirms that the codebook-guided summaries provide cognitive gains for human participants in perceiving places of interest (POIs)’ social inclusiveness. Our study marks an important step toward understanding implicit social barriers and inequalities, demonstrating the great potential of promoting social inclusiveness with Web technology. [^1] This paper has been accepted at The ACM Web Conference 2026. Please cite the version appearing in the conference proceedings.
Segregation; Social Media; Large Language Model; Human Mobility
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3793020ccs: Applied computing Sociologyccs: Computing methodologies Information extractionccs: Human-centered computing Collaborative and social computing
1. Introduction
@@
The rapid advancement of web technologies has enabled fine-grained records of social interactions through diverse media, such as textual posts (Weller et al., 2014), photos (Gilbert et al., 2013), and short videos (Chen et al., 2024). As a result, web platforms have become extensive social sensing systems, generating rich digital traces that reflect social dynamics in the physical world (Stier et al., 2021; Iqbal et al., 2023). Such data offers a unique opportunity to uncover subtle patterns or preferences in individuals’ everyday life, allowing researchers to study long-standing social issues from a new perspective (Moro et al., 2021). Among the most enduring of these issues is segregation: the systematic separation of individuals or groups based on certain characteristics. However, identifying cues of segregation experience from massive, multimodal web content poses significant challenges, requiring advanced methods to capture nuanced feelings, and there is no established procedure to guide this process.
Historically, research on segregation has focused primarily on residential patterns (Park and Burgess, 1925), where individuals of the same racial or ethnic group are more likely to reside in the same neighborhoods (Schelling, 2006). This type of spatial segregation has been linked to negative social outcomes, such as limited upward social mobility and increased crime rates (Gordon and Monastiriotis, 2006; Kramer and Hogue, 2009). Recent studies have shifted focus to experienced segregation—the dynamic segregation individuals experience in daily movements (Moro et al., 2021). Despite the seemingly free human movements in most modern urban spaces, certain demographic groups continue to face barriers to equal access and social interactions. In other words, there seem to be “invisible walls” in cities that prevent sufficient social interactions between groups. These walls are not merely the result of physical proximity but are influenced by cultural, social, and economic factors that drive segregation on a more nuanced level. Accurately predicting segregation experiences can inform individuals to avoid potentially uncomfortable situations and support policymakers in fostering social inclusiveness. Moreover, research suggests that segregation experiences can intensify in larger urban areas (Nilforoshan et al., 2023), underscoring the need to explore the underlying mechanisms for growing more equitable cities admist ever-increasing urbanization. However, existing studies only present retrospective and holistic measurements, providing limited insights and prediction capabilities.
@@
To address these challenges, we propose to leverage the reasoning power of LLMs for automated social media content analysis. We propose a reflective LLM coder, featuring a strategic agentic workflow that integrates two key components: a reflective attributor and a code summarizer. The reflective attributor employs an abductive reasoning approach: it prompts the LLM to estimate a location’s appeal to different demographic groups from its social media content, and reconcile discrepancies through iterative reflection on observed segregation patterns. Subsequently, the code summarizer applies chain-of-thought reasoning (Wei et al., 2022) to iteratively merge insights into a structured codebook. This codebook guides LLMs to transform free-text reviews into structured summaries, highlighting segregation-related factors like cultural resonance and community engagement. Complementing this qualitative approach, we propose a REasoning and EMbedding (RE’EM) framework for quantitative segregation prediction, which combines LLMs’ reasoning capabilities with the representational power of pre-trained embedding models. First, we prompt the LLM to provide structured ratings of a place’s appeal to different groups based on the codebook, allowing the reasoning outcomes of LLMs to be vectorized and easily integrated with other channels of features. Concurrently, we finetune an embedding model to learn global representations optimized for segregation prediction. Finally, we fuse structured ratings, global embeddings, and population information using a neighbor-aware multi-view predictor.
@@
2. Related Work
@@
Mining Web Data with LLMs. The digital footprints and rich textual contents people generate on Web platforms have proven to be informative for a wide array of social phenomena, including health outcome (Nguyen et al., 2016), mental well-being (Mitchell et al., 2013), crime rates (Fatehkia et al., 2019), and neighborhood disparities (Rama et al., 2020; Iqbal et al., 2023). However, these studies often rely on pre-calculated indices or pre-defined word lists to engineer and extract features. This is not only labor-intensive but also lacks adaptability to evolving research questions and contexts. Given the remarkable language understanding and reasoning capabilities of LLMs, recent research explored the potential of leveraging LLMs to mine Web data for social good, such as revealing food-related social prejudice (Luo et al., 2024), evaluating public accessibility (Li et al., 2024), and capturing urban perception (Santos et al., 2024). Nevertheless, these studies often lack an examined framework for insight extraction from massive social media content, which limits their effectiveness in capturing the complexity of Web data. In contrast, our work presents the first step toward unlocking the reasoning power of LLMs to code free-form online reviews, extracting human-comprehensible, informative insights for segregation prediction.
Experienced Segregation. The study of segregation traces its roots back to the early 20th century, when sociologists examined the division of urban spaces into “ecological niches”, each inhabited by distinct social groups (Park and Burgess, 1925). Researchers revealed the detrimental effects of enduring segregation even in the absence of “legalized racial segregation”, on education, income, housing, and crime (King and Mieszkowski, 1973; Massey et al., 1987; Charles, 2003). Albeit offering valuable insights, these studies are highly constrained by data availability, overlooking the reality that individuals spend significant time and engage in numerous interactions beyond the confines of their residential neighborhoods (Wang et al., 2018). With the rapid proliferation of smart mobile devices, the ability to track people’s intricate movements in urban spaces has emerged (Xu et al., 2016). This has paved the way for a novel research avenue aimed at quantifying and understanding the type of segregation individuals actually experience in their daily movements (Moro et al., 2021; Athey et al., 2021; de la Prada and Small, 2024), including how it may change during disasters (Yabe et al., 2023; Chen et al., 2023) and scale with city sizes (Nilforoshan et al., 2023). Nevertheless, existing studies primarily focus on static spatial distribution of demographic groups and physical movement, overlooking the complex socioeconomic factors behind the segregation phenomenon, such as cultural resonance and local involvement. Thus, these works can only provide retrospective measurement but have limited predictive power. In contrast, our work establishes an LLM-based method to automatically extract nuanced features from online reviews for experienced segregation prediction.
Figure 1. Overview of the reflective LLM coder. It consists of a reflective attributor that integrates multi-source review and image signals and refines insights using real visitation patterns, and a code summarizer that consolidates these insights into a structured, generalizable codebook capturing key factors shaping experienced segregation.
3. Preliminary
3.1. Problem Formulation
@@
Inspired by (Moro et al., 2021), we compute the proportion of visitors at POI $i$ by each racial group $q$, denoted as $\tau_{qi}$. We then quantify the segregation at POI $i$ as the deviation of the visitor proportion $\tau_{qi}$ from the city’s residential proportion $T_{q}$ (k is a constant that normalizes $S_{i}$ to range between 0 and 1):
(1) $S_{i}=k\sum_{q}(|\tau_{iq}-T_{q}|).$
@@
(2) $min~\text{diff}(f(i,c,p),S_{i}).$
3.2. Data
@@
Mobility data. To derive the ground truth of experienced segregation, we utilize the Safegraph Patterns Dataset (accessed through Advan) [^4] https://www.deweydata.io/data-partners/advan. This dataset records monthly visits to each POI originating from various CBGs. We aggregate one year’s visitation records to construct a robust dataset with extensive observations. To address the distinct POI indexing systems between Yelp and SafeGraph, we match POIs by ensuring identical names and a location deviation within 200 meters. In Appendix A, we show the variability of POI features, highlighting the importance of our study.
We use demographic and mobility data from 2019 for temporal consistency. To leverage historical POI information, we include social media content from 2019 (16%) and the preceding eight years (84%). Each review is timestamped, enabling LLM to reason about the temporal relationship between the content and the current analysis period.
4. Reflective LLM Coder
The vast volume of raw social media content creates a significant cognitive barrier to revealing the encoded segregation experience. Therefore, an effective approach is needed to process and extract insights. Building on research demonstrating the consistency and stability of LLM-based coding (Dai et al., 2023; Tai et al., 2024), we design a reflective LLM coder to generate an insightful codebook from a small set of POI data, capturing the multi-faceted factors influencing segregation and offering a structured framework for deeper analysis. As shown in Figure 1, it contains two modules, reflective attributor and code summarizer.
4.1. Reflective Attributor
We design a reflective attributor to identify factors influencing a POI’s appeal to different groups. Specifically, we design a two-step Chain-of-Thought (CoT) scheme (Wei et al., 2022) that guides an LLM to integrate multiple perspectives from texts and images, forming a reflection-driven process that refines insights based on real-world feedback.
@@ 4.1. Reflective Attributor
Table 1. Automatically constructed codebook. Each dimension represents a distinct driver of POI attractiveness or deterrence for different demographic groups, forming the conceptual “bricks” underlying experienced segregation patterns. The explanations summarize the thematic meaning of each code.
@@
Reflection-guided insight refinement. Importantly, we introduce real-world data to guide LLM’s reflections upon these insights. We measure the difference between the demographic composition of visitors to each POI and that of nearby residents, using it as a ground-truth signal to identify which racial/ethnic groups it attracts or repels. This signal guides the LLM in refining the existing insights in an abductive manner, retaining only those consistent with real-world evidence. This step not only enhances the relevance of the insights, but also provides a grounded explanation for observed visitation preferences.
Figure 2. Reasoning-and-Embedding (RE’EM) framework. RE’EM integrates three complementary channels (reasoning, embedding, and population) with a neighbor-aware multi-view fusion to predict POIs’ experienced segregation.
@@ 4.2. Code Summarizer
5. REasoning-and-EMbedding (RE’EM) Framework
@@ 5.1. Reasoning Channel
(3) $r_{i},e_{i}=\text{LLM}(u_{i},\text{codebook}).$
@@
(4) $v_{i}^{r}=\text{ReasonAdapter}(r_{i}).$
5.2. Embedding Channel
The embedding channel extracts deep semantic representations $v_{i}^{e}$ from the review corpus associated with each POI $i$. Due to the token length constraints, we apply data augmentation, i.e., sample different subsets of $c_{i}$ for input to the embedding model. We fine-tune an open-source text embedding model, GTE-base (Li et al., 2023), to enhance its capability of extracting higher-level representations while preserving the pre-trained semantic knowledge. Although both the reasoning and embedding channels take $c_{i}$ as input, they extract complementary aspects: structured attribute reasoning in the former, and holistic semantic embedding in the latter. The resulting embedding is then passed through an embedding adapter (MLP) to produce $v_{i}^{e}$.
(5) $v_{i}^{e}=\text{EmbeddingAdapter}(\text{GTE}(c_{i})).$
5.3. Population Channel
@@
(6) $v_{i}^{p}=\text{PopulationAdapter}(p_{i}).$
5.4. Neighbor-Aware Multi-View Predictor
@@
(8) $\hat{S_{i}}=\text{Fuse}(v_{i}^{r},v_{n_{i}}^{r},v_{i}^{e},v_{n_{i}}^{e},v_{i}^{p},v_{n_{i}}^{p}).$
Table 2. Segregation prediction performances across different method. Compared with population-only and embedding baselines, RE’EM achieves significant improvements across MSE, RMSE, MAE, and R².
@@
Qwen3-Embedding-8B 0.0072±0.0002 4.00% (p=0.0196) 0.0848±0.0011 1.85% (p=0.0292) 0.0683±0.0009 0.00% (p=0.4789) 0.3349±0.0181 5.85% (p=0.0812)
RE‘EM 0.0068±0.0003 9.33% (p=0.0080) 0.0823±0.0021 4.75% (p=0.0071) 0.0662±0.0021 3.07% (p=0.0597) 0.3885±0.0278 22.79% (p=0.0022)
6. User Study
We conduct a user study to evaluate the cognitive gain to humans provided by our generated codebook. We recruit participants via snowball sampling among graduate students and researchers with hands-on experience or sufficient knowledge in social media studies: approximately 17.3% from urban planning, 24% from social science, and 58.7% from data mining. To ensure that participants fully understood the survey tasks, we provided clear explanations of the geographic and racial concepts required for the study. Our questionnaire is included in our released code repository  [^5] https://github.com/tsinghua-fib-lab/Invisible-Wall-in-Cities.